OpenStudy (anonymous):

Sorry thr question reads -6/7 x (1/6)=

OpenStudy (anonymous):

\[\LARGE -\frac67 \times \frac16\] ? do you know to multiply fractions? \[\LARGE \frac ab \times \frac cd =\frac{ac}{bd}\] does that help? :)

OpenStudy (anonymous):

So is it -1/7